# Catalytic Quantum Error Correction: Theory, Efficient Catalyst Preparation,and Numerical Benchmarks

### 31 Mar 2026 v2

@@ I Introduction
Moreover, even assuming the theoretical validity of catalytic recovery, the practical bottleneck is catalyst preparation. The constructive proof of Ref. [31] requires $n^{*}\sim d^{4}e^{2\gamma}$ noisy copies to prepare a catalyst with fidelity $F_{\mathrm{cat}}\geq 0.99$, where $d$ is the Hilbert space dimension and $\gamma$ is the dephasing strength. For a 6-qubit system ($d=64$) under moderate dephasing ($\gamma=2$), this translates to $n^{*}\approx 7.3\times 10^{10}$ copies—far beyond current experimental capabilities. Recent advances in quantum state purification [8, 41, 42] and dynamical decoupling [37, 26] offer paths to dramatically better scaling.
@@ II.2 Coherent modes and mode inclusion
Crucially, the mode inclusion condition is both necessary and sufficient: if $\mathcal{C}(\rho^{\prime})\not\subseteq\mathcal{C}(\rho)$, then $R(\rho\to\rho^{\prime})=0$ and no catalytic transformation can achieve the conversion [31]. This provides a sharp classification of state pairs into recoverable ($R=\infty$) and irrecoverable ($R=0$), with no intermediate regime.
@@ II.3 CQEC protocol
The CQEC protocol applies Theorem 2 to quantum state recovery. We use the term “recovery” rather than “correction” to emphasize the key operational distinction: CQEC recovers a known target state from noisy copies, whereas QEC corrects errors on an unknown encoded state. The requirement that $\rho_{0}$ be known does not reduce CQEC to trivial state re-preparation: the catalyst enables recovery from noisy copies without re-executing the (potentially expensive) state preparation circuit, provided sufficient copies are available. CQEC does not violate the no-cloning theorem [40] for three reasons: (i) the target state $\rho_{0}$ is known, so no unknown-state information is being duplicated; (ii) the $n$ input copies are consumed (not preserved) by the protocol; and (iii) the covariant channel $\Lambda$ is state-specific and cannot be applied to arbitrary unknown states.
@@
where $n_{\mathrm{cat}}$ is the number of catalytic branches, $\tau_{n_{\mathrm{cat}}-k}=\mathrm{Tr}_{1,\ldots,k}[\sigma_{n_{\mathrm{cat}}}]$ is the reduced state, and $R$ is a register system of dimension $n_{\mathrm{cat}}$. The apparent circularity is resolved by the constructive proof of Ref. [31]: two-level catalysts $c_{i}$ are first constructed from $\mu_{i}$ copies via non-catalytic covariant operations, then composed into the full catalyst via Theorem S.10. The catalyst dimension is $\dim(c)=n_{\mathrm{cat}}\cdot d^{n_{\mathrm{cat}}-1}\cdot n_{\mathrm{cat}}$, which grows exponentially. In our variational implementation (Sec. III.3), we use a simplified catalyst of dimension $d_{C}=2^{n_{C}}$ with $n_{C}=n_{S}$ qubits. This gap is fundamental: the variational circuit searches for a high-fidelity covariant transformation within a restricted ansatz, validating theoretical predictions but not implementing the full constructive proof. Efficient catalyst preparation strategies that dramatically reduce the required resources are developed in Sec. IV.
@@ III.4 Parameter optimization
Approximation quality vs. dimension. The fidelity gap grows with $d$: $1-F<10^{-8}$ for $d=4$, $1-F<10^{-4}$ for $d=8$, $1-F\approx 10^{-4}$ for $d=16$, and $1-F\approx 2\times 10^{-4}$ for $d=64$. Increasing to 10–15 gates recovers $10^{-6}$ for all $d$. The covariant unitary group for $n=8$ qubits has $\sum_{k}\binom{n}{k}^{2}-1=12\,869$ free parameters [30]; our 5–15 parameter ansatz explores a small submanifold, with high fidelity implying the target $\Lambda$ lies near this submanifold.
@@ IV.3 Recursive swap test purification
The swap test [8, 1] on two copies $\rho$ and $\sigma$ projects onto the symmetric subspace $\mathrm{Sym}^{2}(\mathbb{C}^{d})$ via the projector $\Pi_{\mathrm{sym}}=(I+\mathrm{SWAP})/2$:
@@ VI.1 CQEC asymptotic benchmark
Table 7: CQEC recovery fidelity (mean $\pm$ std, 10 seeds). Success rate is 100% for all entries. Regev results (below rule) serve as a high-dimensional stress test; CQEC is inapplicable to factoring in practice due to the target-state requirement (see Sec. VII.3).
@@ VI.10 Resource overhead
Table 11: Resource comparison for $n_{\mathrm{logical}}$ logical qubits. CQEC qubit count is $n_{S}+n_{C}+n_{A}+n_{\mathrm{copies}}$; gate count is for the variational recovery circuit only and does not include the cost of preparing $n^{*}$ noisy copies (Table 9). With DD+Twirl, $n^{*}=8$3232, adding at most $32\times G_{\mathrm{alg}}$ gates.
@@
1 8 / 5 7 / 10 13 / 36 41 / 100
3 22 / 11 21 / 30 39 / 108 123 / 300
6 137 / 20 42 / 60 78 / 216 246 / 600
The total CQEC resource cost includes: (i) preparation of $n$ noisy copies ($n\times G_{\mathrm{alg}}$ gates), (ii) DD pulses ($n\times N$ single-qubit gates), (iii) Clifford twirling ($n\times O(n_{q}^{2})$ gates), (iv) $\log_{2}n$ rounds of swap tests ($O(n_{q}\log n)$ two-qubit gates), and (v) the recovery circuit ($G_{\mathrm{CQEC}}=5$1515 EC gates). The total gate count is $n\times(G_{\mathrm{alg}}+N+O(n_{q}^{2}))+O(n_{q}\log n)+G_{\mathrm{CQEC}}$. For qDRIFT with 80 gates and $n^{*}=4.4\times 10^{5}$, the total is $\sim 3.5\times 10^{7}$ gates—more than surface code overhead. CQEC’s advantage is therefore not in absolute resource cost but in threshold-free recovery. For systems relevant to quantum advantage ($n\geq 50$), the CQEC copy overhead scales as $\mu\sim 2^{2n}e^{2\gamma}$, which is exponential—far exceeding the polynomial overhead of surface codes. A practically relevant regime for CQEC may be small quantum modules ($n\leq 10$) within a larger fault-tolerant architecture.
@@ VII Discussion
VII.1 Two paradigms for quantum information protection
Our results reveal a complementary relationship between two protection strategies:
@@ VII.2 Complementarity of catalyst preparation methods
VII.3 Limitations
@@
VII.4 Mode no-broadcasting and fundamental limits
@@ VIII Conclusion
The main practical contribution is the DD+Twirl+Swap Test pipeline for catalyst preparation. CPMG dynamical decoupling reduces effective dephasing ($\gamma_{\mathrm{eff}}=\gamma/(N+1)$), Clifford twirling converts residual anisotropic dephasing into isotropic depolarizing noise, and the standard swap test achieves doubly exponential purification. With CPMG-8 and 8 copies, the pipeline achieves $F_{\mathrm{cat}}>0.96$ uniformly across d=4d=46464 under strong dephasing ($\gamma=2$)—a $10^{9}$-fold improvement over distillation. The total overhead is 8 DD pulses per copy plus Clifford twirling gates, negligible compared to the copy-count savings.
Our results suggest that coherence resource theory offers a promising direction for quantum state recovery that is fundamentally distinct from the stabilizer formalism, though substantial challenges remain. Three concrete next steps toward practical implementation are: (i) experimental demonstration of the DD+Twirl pipeline on superconducting or trapped-ion platforms for $d\leq 16$, where the copy overhead (n=8n=83232) is within reach; (ii) proving tight analytical bounds on the pipeline’s fidelity as a function of DD pulse count $N$, twirl sample count $K$, and copy number $n$, closing the gap between our numerical observations and rigorous guarantees; and (iii) integration with fault-tolerant architectures, where surface codes protect data qubits while CQEC restores coherence in small ancilla registers ($d\leq 32$) used for state preparation subroutines.

### 06 Apr 2026 v3

@@ I Introduction
Concurrently with this work, Raghoonanan and Byrnes [29] introduced Purification QEC (PQEC), which also uses $N$ noisy copies and recursive swap tests to recover quantum states, but does not require knowledge of the target state. PQEC achieves doubly exponential purification ($\rho\mapsto\rho^{N}/\mathrm{Tr}(\rho^{N})$) with $O(M\log_{2}N)$ qubits, and tolerates depolarizing errors up to $p=3/4$. The relationship between CQEC and PQEC illuminates a fundamental trade-off: PQEC is state-agnostic but bounded by the noise threshold ($p_{\mathrm{th}}=3/4$ depolarizing, $1/2$ dephasing), while CQEC exploits target-state knowledge to achieve threshold-free recovery at the cost of requiring this additional information. We provide a detailed comparison in Sec. VII.3.
Moreover, even assuming the theoretical validity of catalytic recovery, the practical bottleneck is catalyst preparation. The constructive proof of Ref. [32] requires $n^{*}\sim d^{4}e^{2\gamma}$ noisy copies to prepare a catalyst with fidelity $F_{\mathrm{cat}}\geq 0.99$, where $d$ is the Hilbert space dimension and $\gamma$ is the dephasing strength. For a 6-qubit system ($d=64$) under moderate dephasing ($\gamma=2$), this translates to $n^{*}\approx 7.3\times 10^{10}$ copies—far beyond current experimental capabilities. Recent advances in quantum state purification—including the recursive swap test of Childs et al. [8] (which we adopt in Sec. IV.3) and related protocols [42, 43]—together with dynamical decoupling [38, 26] offer paths to dramatically better scaling.
@@ II.2 Coherent modes and mode inclusion
Crucially, for full-rank $\rho$, the mode inclusion condition is both necessary and sufficient: if $\mathcal{C}(\rho^{\prime})\not\subseteq\mathcal{C}(\rho)$, then $R(\rho\to\rho^{\prime})=0$ and no catalytic transformation can achieve the conversion [32]. (The full-rank requirement is satisfied automatically for any finite noise channel applied to a pure state.) This provides a sharp classification of state pairs into recoverable ($R=\infty$) and irrecoverable ($R=0$), with no intermediate regime.
@@ II.3 CQEC protocol
The CQEC protocol applies Theorem 2 to quantum state recovery. We use the term “recovery” rather than “correction” to emphasize the key operational distinction: CQEC recovers a known target state from noisy copies, whereas QEC corrects errors on an unknown encoded state. Despite the name “error correction,” CQEC is more precisely a catalytic state recovery protocol. We retain “error correction” in the name because (i) the protocol corrects the effects of decoherence on quantum states, and (ii) the catalyst-mediated recovery is reusable across cycles, analogous to syndrome-based correction rounds. The requirement that $\rho_{0}$ be known does not reduce CQEC to trivial state re-preparation or tomography: the covariant channel $\Lambda$ operates entirely in the quantum domain without measurement, preserving entanglement and coherence structure that tomography would destroy. The practical value lies in recovering expensive-to-prepare states cheaply from noisy copies rather than re-executing the original circuit. CQEC does not violate the no-cloning theorem [41] for three reasons: (i) the target state $\rho_{0}$ is known, so no unknown-state information is being duplicated; (ii) the $n$ input copies are consumed (not preserved) by the protocol; and (iii) the covariant channel $\Lambda$ is state-specific and cannot be applied to arbitrary unknown states.
@@
where $n_{\mathrm{cat}}$ is the number of catalytic branches, $\tau_{n_{\mathrm{cat}}-k}=\mathrm{Tr}_{1,\ldots,k}[\sigma_{n_{\mathrm{cat}}}]$ is the reduced state, and $R$ is a register system of dimension $n_{\mathrm{cat}}$. The apparent circularity is resolved by the constructive proof of Ref. [32]: two-level catalysts $c_{i}$ are first constructed from $\mu_{i}$ copies via non-catalytic covariant operations, then composed into the full catalyst via Theorem S.10. The catalyst dimension is $\dim(c)=n_{\mathrm{cat}}\cdot d^{n_{\mathrm{cat}}-1}\cdot n_{\mathrm{cat}}$, which grows exponentially. In our variational implementation (Sec. III.3), we use a simplified catalyst of dimension $d_{C}=2^{n_{C}}$ with $n_{C}=n_{S}$ qubits. This gap is fundamental: the variational circuit searches for a high-fidelity covariant transformation within a restricted ansatz, validating theoretical predictions but not implementing the full constructive proof. Efficient catalyst preparation strategies that dramatically reduce the required resources are developed in Sec. IV. Equation (6) defines the formal protocol; in practice, we replace it with the DD+Twirl+Swap Test pipeline (Sec. IV.5) for catalyst preparation and the variational circuit (Sec. III.3) for the recovery step.
@@ III.4 Parameter optimization
Approximation quality vs. dimension. The fidelity gap grows with $d$: $1-F<10^{-8}$ for $d=4$, $1-F<10^{-4}$ for $d=8$, $1-F\approx 10^{-4}$ for $d=16$, and $1-F\approx 2\times 10^{-4}$ for $d=64$. Deeper circuits with more gates are expected to improve fidelity; a systematic study of gate-count scaling is left to future work. The covariant unitary group for $n=8$ qubits has $\sum_{k}\binom{n}{k}^{2}-1=12\,869$ free parameters [31]; our 5–15 parameter ansatz explores a small submanifold, with high fidelity implying the target $\Lambda$ lies near this submanifold.
@@ IV.3 Recursive swap test purification
[…]
@@ VI.1 CQEC asymptotic benchmark
Table 7: CQEC recovery fidelity (mean $\pm$ std, 10 seeds). Success rate is 100% for all entries. Regev results (below rule) serve as a high-dimensional stress test; CQEC is inapplicable to factoring in practice due to the target-state requirement (see Sec. VII.4).
@@ VI.10 Resource overhead
Table 11: Resource comparison for $n_{\mathrm{logical}}$ logical qubits. CQEC qubit count is the recovery circuit only: $n_{S}+n_{C}+n_{A}$ (system + catalyst + ancilla). Noisy copies for catalyst preparation (8–32 with DD+Twirl) are prepared sequentially and do not require simultaneous qubits. Gate count is for the recovery circuit; total cost including copies is $n\times G_{\mathrm{alg}}+G_{\mathrm{CQEC}}$ (see text).
@@
1 4 / 5 7 / 10 13 / 36 41 / 100
3 8 / 11 21 / 30 39 / 108 123 / 300
6 14 / 20 42 / 60 78 / 216 246 / 600
The total CQEC resource cost includes: (i) preparation of $n$ noisy copies ($n\times G_{\mathrm{alg}}$ gates), (ii) DD pulses ($n\times N$ single-qubit gates), (iii) Clifford twirling ($n\times O(n_{q}^{2})$ gates), (iv) $\log_{2}n$ rounds of swap tests ($O(n_{q}\log n)$ two-qubit gates), and (v) the recovery circuit ($G_{\mathrm{CQEC}}=5$1515 EC gates). The total gate count is $n\times(G_{\mathrm{alg}}+N+O(n_{q}^{2}))+O(n_{q}\log n)+G_{\mathrm{CQEC}}$. For qDRIFT with 80 gates and $n^{*}=4.4\times 10^{5}$, the total is $\sim 3.5\times 10^{7}$ gates—more than surface code overhead. CQEC’s advantage is therefore not in absolute resource cost but in threshold-free recovery. For systems relevant to quantum advantage ($n\geq 50$), the CQEC copy overhead scales as $\mu\sim 2^{2n}e^{2\gamma}$, which is exponential—far exceeding the polynomial overhead of surface codes. A practically relevant regime for CQEC may be small quantum modules ($n\leq 10$) within a larger fault-tolerant architecture (Table 11).
@@ VII Discussion
VII.1 Landscape of quantum state protection
Our results, together with concurrent work on PQEC [29], reveal a rich landscape of quantum state protection strategies that differ along three axes: target-state knowledge (known vs. unknown), resource type (redundancy vs. copies vs. catalysts), and noise tolerance (threshold-bounded vs. threshold-free). We highlight two specific complementary relationships:
@@ VII.2 Complementarity of catalyst preparation methods
VII.3 Relation to purification QEC and the state estimation objection
A natural objection is that CQEC, by requiring knowledge of the target state $\rho_{0}$ and multiple noisy copies, is “not really QEC” but rather a form of state estimation or re-preparation. We address this by positioning CQEC within a taxonomy of copy-based quantum state recovery methods, including the concurrent Purification QEC (PQEC) of Raghoonanan and Byrnes [29].
Table 12 compares the three paradigms along seven operational axes.
Table 12: Comparison of quantum state protection paradigms.
Property Conventional QEC PQEC [29] CQEC (this work)
Target state knowledge Not required Not required Required
Input Encoded state + syndromes NN noisy copies $n$ noisy copies + catalyst
Output Corrected state (coherent) Purified state (via $\rho^{N}/\mathrm{Tr}(\rho^{N})$) Recovered state (coherent)
Error threshold Quantitative: $p_{\mathrm{th}}\approx 1\%$ (surface) Quantitative: $p_{\mathrm{th}}=3/4$ (depol), $1/2$ (deph) Qualitative: mode support only
Success condition $p<p_{\mathrm{th}}$ $F>1/D$ (dominant eigenvec. $|\psi\rangle$) Mode inclusion (asymptotic); $n\geq n^{*}$ (finite)
Qubit overhead $O(d^{2})$ per logical qubit $O(M\log_{2}N)$ $(2n_{q}\!+\!2)+2n_{q}\log_{2}n$
Anisotropic noise Corrects via syndromes $p_{\mathrm{th}}^{\mathrm{deph}}=1/2$ (improvable via twirling) Threshold-free (with DD+Twirl)
Interleaving in circuit Native (syndrome cycles) Native (between unitaries) Post-hoc (see text)
CQEC vs. state estimation. State estimation (tomography + re-preparation) destroys the quantum state through measurement, requiring $O(d^{2}/\epsilon^{2})$ measurements and producing a classical description from which a new state is prepared. CQEC, by contrast, operates as a coherent quantum channel during recovery: the covariant map $\Lambda$ transforms $\rho_{\mathrm{noisy}}\otimes c$ to a quantum state $\tau$ whose system marginal approximates $\rho_{0}$, without measurement during the recovery step. (Target-state knowledge is used offline to optimize the circuit parameters $\boldsymbol{\theta}$, not during the recovery channel itself.) The catalyst enables coherent amplification that preserves entanglement structure—a feature impossible via tomography. The “target state knowledge” requirement enters only through the optimization of the circuit parameters $\boldsymbol{\theta}$, not through the recovery channel itself: once $\boldsymbol{\theta}$ is fixed, the circuit acts as a standard quantum channel.
CQEC vs. PQEC. PQEC achieves $\rho^{N}/\mathrm{Tr}(\rho^{N})$ via the recursive swap test, converging to the dominant eigenvector of $\rho$. This is state-agnostic but requires that $|\psi_{0}\rangle$ is the dominant eigenvector—a condition that can fail under anisotropic noise. Under pure dephasing, PQEC’s Bloch vector direction is preserved but the purification map cannot increase the fidelity beyond the direction-dependent ceiling, yielding $p_{\mathrm{th}}^{\mathrm{deph}}=1/2$ [29]. Twirling can boost this threshold toward $3/4$ at the cost of additional Clifford gates [29]. CQEC’s mode inclusion criterion is orthogonal: it asks whether the support of coherence is preserved, not whether the dominant eigenvector aligns with $|\psi_{0}\rangle$. Under partial dephasing ($\gamma<\infty$), all modes survive and CQEC succeeds regardless of $\gamma$ in the asymptotic limit ($n\to\infty$)—even beyond the regime where PQEC saturates without twirling. At finite $n$, however, CQEC’s fidelity gap $1-F\sim C/\sqrt{n}$ grows with $\gamma$ through the constant $C\sim e^{\gamma}$, so very strong dephasing requires proportionally more copies (Table 9). The threshold-free property is thus an asymptotic statement; at finite resources, CQEC and PQEC face different but analogous resource–fidelity trade-offs.
The methods are complementary:
  • PQEC is preferred when the target state is unknown and noise is predominantly depolarizing ($p<3/4$).
  • CQEC is preferred when the target state is known and noise is anisotropic (dephasing-dominated), where PQEC’s threshold limits recovery.
  • A hybrid approach is natural: PQEC first purifies $N$ noisy copies state-agnostically, producing a single copy with $F\approx 1-\epsilon$; this higher-fidelity copy then serves as CQEC’s input, where mode inclusion (preserved by PQEC’s spectral map) enables threshold-free final recovery. The interface is straightforward because PQEC’s output is a valid density matrix that can be fed directly into the CQEC covariant channel.
Interleaving limitation. Unlike PQEC, which can be interleaved between unitary steps of a quantum algorithm (because the swap test commutes with bilateral rotations $\Pi_{\pm},U\otimes U]=0 [$29]), CQEC operates post-hoc on the algorithm’s output state. This is because the covariant recovery channel $\Lambda$ is state-specific: the optimized parameters $\boldsymbol{\theta}$ depend on the target state $\rho_{0}$, which changes at each intermediate step of the algorithm. Interleaving CQEC would require re-optimizing $\boldsymbol{\theta}$ at every step—computationally expensive but not fundamentally impossible. In a hybrid architecture, PQEC could interleave purification during the algorithm while CQEC provides final state-specific recovery at the output.
CQEC’s unique advantage. CQEC’s distinctive contribution is the catalytic covariant recovery channel $\Lambda$, which achieves threshold-free recovery in the asymptotic limit—a property no state-agnostic method (including PQEC) can match. PQEC converges to the dominant eigenvector of $\rho$; when this eigenvector does not coincide with $|\psi_{0}\rangle$ (as under strong dephasing without twirling), PQEC’s fidelity saturates below unity regardless of copy count [29]. CQEC, by exploiting knowledge of $\rho_{0}$ to construct a state-specific $\Lambda$, overcomes this limitation. The DD+Twirl pipeline is a general noise-preprocessing technique applicable to any copy-based protocol; its development here was motivated by CQEC’s catalyst preparation but could benefit PQEC as well.
VII.4 Limitations
@@
VII.5 Mode no-broadcasting and fundamental limits
@@ VIII Conclusion
The paper makes two logically independent contributions: (1) the CQEC recovery protocol itself (covariant channel $\Lambda$ mediated by a catalyst), and (2) the DD+Twirl+Swap Test pipeline for efficient catalyst preparation. The pipeline (contribution 2) is a general-purpose noise preprocessing technique equally applicable to other copy-based protocols including PQEC [29]. The CQEC-specific contribution (1) is the threshold-free recovery via catalytic covariance. CPMG dynamical decoupling reduces effective dephasing ($\gamma_{\mathrm{eff}}=\gamma/(N+1)$), Clifford twirling converts residual anisotropic dephasing into isotropic depolarizing noise, and the standard swap test achieves doubly exponential purification. With CPMG-8 and 8 copies, the pipeline achieves $F_{\mathrm{cat}}>0.96$ uniformly across d=4d=46464 under strong dephasing ($\gamma=2$)—a $10^{9}$-fold improvement over distillation. The total overhead is 8 DD pulses per copy plus Clifford twirling gates, negligible compared to the copy-count savings.
Our results suggest that coherence resource theory offers a promising direction for quantum state recovery that is fundamentally distinct from both the stabilizer formalism and state-agnostic purification methods such as PQEC [29]. CQEC and PQEC occupy complementary niches: PQEC operates without target-state knowledge but is bounded by noise thresholds ($p_{\mathrm{th}}=3/4$ depolarizing, $1/2$ dephasing), while CQEC exploits target-state knowledge to achieve threshold-free recovery under anisotropic noise. Substantial challenges remain. Three concrete next steps toward practical implementation are: (i) experimental demonstration of the DD+Twirl pipeline on superconducting or trapped-ion platforms for $d\leq 16$, where the copy overhead (n=8n=83232) is within reach; (ii) proving tight analytical bounds on the pipeline’s fidelity as a function of DD pulse count $N$, twirl sample count $K$, and copy number $n$, closing the gap between our numerical observations and rigorous guarantees; and (iii) integration with fault-tolerant architectures, where surface codes protect data qubits while CQEC restores coherence in small ancilla registers ($d\leq 32$) used for state preparation subroutines.